7) The path of a falling object is given by the function where represents the initial velocity in ft/sec and represents the initial height in feet.
a) If a rock is thrown upward with an initial velocity of 64 feet per second from the top of a 25-foot building, write the height (s) equation using this information.
Typing hint: Type t-squared as t^2
Answer:
b) How high is the rock after 1 second?
Answer:
Show your work here:

c) After how many seconds will the graph reach maximum height?
Answer:
Show your work here:

d) What is the maximum height?
Answer:
Show your work here:

Answer by Alan3354(69443)

a) If a rock is thrown upward with an initial velocity of 64 feet per second from the top of a 25-foot building, write the height (s) equation using this information.
Typing hint: Type t-squared as t^2
Answer: h(t) = -16t^2 + 64t + 25 (height above the ground, not the building)
---------------
b) How high is the rock after 1 second?
Answer: h(1) = -16 + 64 + 25 = 73 feet
Show your work here:
-------------------------
c) After how many seconds will the graph reach maximum height?
Answer: 2
Show your work here: Solve for h = 25 feet.
-16t^2 + 64t + 25 = 25
16t^2 = 64t
t = 0, t = 4
The rock is at 25 feet twice, once when launched (t=0) and on its way down (t=4). The time at max height is the center of the 2 times, = 2 seconds.
d) What is the maximum height?
Answer: h(2) = -64 + 128 + 25 = 89 feet
Show your work here: